OpenStudy (anonymous):

OMG...the table lists data regarding the average salaries of several proffesional athletes in the years 1991 and 2001. a) use the data points to find a linear function that fits the data. b) use the function to predict the avg salary in 2005 and 2010. The predicted avg salary for 2005 is? The predicted avg salary for 2010 is?

OpenStudy (saifoo.khan):

i already answered! :(

OpenStudy (anonymous):

where babe?

OpenStudy (saifoo.khan):

S(x) = 272000+1450,000(14) 20572000 S(x) = 272000+1450,000(2010-1991) = 27822,000

OpenStudy (anonymous):

okay, well I think I got it, but not sure.

OpenStudy (anonymous):

Okay, got a question....

OpenStudy (saifoo.khan):

THANK GOD. lol.

OpenStudy (anonymous):

LOL..now don't be mean. LOL How do I find the slope of a line?

OpenStudy (anonymous):

That just boggles my mind.

OpenStudy (anonymous):

brb///gotta pee again...

OpenStudy (saifoo.khan):

Slope: 1450,000 haha, okayy!

OpenStudy (saifoo.khan):

still peeing?

OpenStudy (anonymous):

no..LOL Iam back.

OpenStudy (saifoo.khan):

lol, i think u were peeing for the whole next year, jk/

OpenStudy (anonymous):

I breed Persian cats and I had to clean their litter boxes.

OpenStudy (anonymous):

LOL....

OpenStudy (saifoo.khan):

OIC,

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

Wasn't peeing all that time.

OpenStudy (saifoo.khan):

haha,ok..

OpenStudy (anonymous):

I am laughing.

OpenStudy (anonymous):

Okay, so you ready to help me some more? I have to have my test in by midnight.

OpenStudy (saifoo.khan):

OH YEs, im ready!

OpenStudy (anonymous):

I am so appreciative of the help you have given me.

OpenStudy (saifoo.khan):

well thnks for that! :)

OpenStudy (anonymous):

okay, lets go....